Investors do not invest all their money in the investment predicted to be the best because it is too risky. An investor holds shares in three independent companies:Company A, Company B, and Company C. An analyst predicts the probability of profit increase for each as follows:
Company A:0.70
Company B:0.50
Company C:0.065
A.)If the analyst's predictions are true, what is the probability that all three companies show a profit increase?
B.)If the analyst's predicitions are true, what is the probability that at least one company will show profit increase?

Answer by stanbon(75887)

An investor holds shares in three independent companies:Company A, Company B, and Company C. An analyst predicts the probability of profit increase for each as follows:
Company A:0.70
Company B:0.50
Company C:0.065
A.)If the analyst's predictions are true, what is the probability that all three companies show a profit increase?
---
P(all show profit) = 0.7*0.5*0.065 = 0.02275...
-----------------------------------------------------
B.)If the analyst's predicitions are true, what is the probability that at least one company will show profit increase?
-----
P(none show profit) = 0.3*0.5*0.935 = 0.14025..
----
P(at least one company shows profit) = 1-P(all show)-P(none show)
= 1 - 0.02275 - 0.14025 = 0.8370...
Cheers,
Stan H.